OpenStudy (anonymous):

which number comes between 0.5 and 8/9 22/24 25/36 7/24 1/3

OpenStudy (aum):

Are you allowed to use a calculator?

OpenStudy (anonymous):

and yes

OpenStudy (aum):

0.5 8/9 = 0.8889 22/24 = 0.9167 25/36 = 0.6944 7/24 = 0.2917 1/3 = 0.3333 Which of the bottom four decimals fall between 0.5 and 0.8889?

OpenStudy (anonymous):

b

OpenStudy (aum):

correct. b) 25/36

OpenStudy (anonymous):

shanks can you help with some more

OpenStudy (aum):

one more and I have to leave after that.

OpenStudy (anonymous):

ok

OpenStudy (aum):

post the question. I am running late.

OpenStudy (anonymous):

well you can go its ok

OpenStudy (aum):

alright.